Some help in this question, please?!
In the casino game of roulette, you can bet on whether the ball will fall in an arc on the wheel colored red, black, or green. The arcs are of equal length with 18 colored red, 18 colored black and 2 colored green. Suppose you make a S5 bet on red. 3. a. Create a probability distribution for this situation where X = net winnings for this single bet. (4 points) b. Compute the expected or mean net winnings for this single bet. Use the correct notation
Answer #1

a)

P(winning)=18/38(red outcome)

P(losing)=20/38(black or green outcome)

If you win, you get $5

If you lose, you lose $5

Let X be the net winnings on single bet

Probability distribution is given by

Earnings,x Probability, P(x)
$5 18/38
-$5 20/38

b)

E(X)=\sum xP(x)

E(X)=(5*\frac{18}{38})+(-5*\frac{20}{38})

E(X)=-0.2632

You lose about 26 cents each time you play this game on an average